OpenStudy (anonymous):

if a+b=4 then, a^3+b^3=?

OpenStudy (anonymous):

a+b=4 Make this one... a = -b + 4 a^3+b^3 Substitute the first equation into the second one (-b + 4)^3 + b^3 Does this help?

OpenStudy (anonymous):

Having this (-b + 4)^3 + b^3 you can now solve for b.

OpenStudy (anonymous):

and then slove for a once u get b right?

OpenStudy (anonymous):

Yepp by plugging b back into the first equation. (a+b=4)

OpenStudy (anonymous):

oh okay thanks :D

geerky42 (geerky42):

"Having this (-b + 4)^3 + b^3 you can now solve for b." But we don't know what a³ + b³ equal to..?

OpenStudy (anonymous):

@geerky42 You have a point here...

OpenStudy (anonymous):

But it also looks like an if-then statement. Now that I look at it,

OpenStudy (anonymous):

Well when I do (-b + 4)^3 + b^3 -b^3+4^3+b^3 b cancels out and it just sits there with 4^3 which is 64 so my way is for sure not the way to do it, if you try a this same way a will cancel out also.

geerky42 (geerky42):

But (-b + 4)³ \(\neq\) -b³ + 4³, it's actually -b³ + 12b² - 48b + 64

OpenStudy (anonymous):

You can't do it... Actually it depends on what a and b equal. Mathematically: (a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3 = (a^3 + b^3) + 3ab(a+b) therefore by rearranging a^3 + b^3 = 3ab(a+b) - (a+b)^3 you have (a+b), but you are missing ab (which would mean you could calculate a and b) If you don't get what I'm talking about, try an example: Let a = 1, b = 3 a + b = 4, a^3+b^3 = 28 Now, Let a = 0, b = 4 a + b = 4, a^3 + b^3 = 64...a different answer. So you can see that a^3 + b^3 depends entirely on what a and b are. You don't have enough information to calculate this

geerky42 (geerky42):